adam is 20 years younger than brian .in 2 years brian willbe as old as adam.how are they old now find the equation

Answer by harpazo(655)

Let A = Adam
Let B = Brian
A = x - 20
B = x
In two years, A = (x-20) + 2 and B = x + 2
Can you take it from here?